OpenStudy (anonymous):

Help pleasee ill give a medal and ill fan you - i finished the first part but im confused about the others

jimthompson5910 (jim_thompson5910):

So you're stuck on this part? "Graph the price the coffee shop would pay for Coffee Delivery B to deliver the same amount of cups on the same graph from Part C, using a different color for the new line."

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

unfortunately, the equation b=3.5x+3.9(500)c is incorrect

jimthompson5910 (jim_thompson5910):

let's make x = number of cups y = total cost

jimthompson5910 (jim_thompson5910):

if they charge 3.9 cents per cup, ie $0.39 per cup, then the total cost per week is y = 0.39x + 3.5

jimthompson5910 (jim_thompson5910):

To graph this, we just need two points Plug in x = 0 y = 0.39x + 3.5 y = 0.39*0 + 3.5 y = 3.5 So one point is (0,3.5) Plug in x = 1 y = 0.39x + 3.5 y = 0.39*1 + 3.5 y = 3.89 Another point is (1,3.89)

jimthompson5910 (jim_thompson5910):

Plot those two points. Then connect them with a straight line.

OpenStudy (anonymous):

that makes sense thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome